# Forecasting When to Forecast: Accelerating Diffusion Models with Confidence-Gated Taylor

### v2

@@ 1 Introduction
In recent years, diffusion models  [6, 29, 31] have become foundational to visual generation, owing to their impressive ability to produce high-quality and semantically rich images. Initially built upon U-Net [26] architectures, these models have demonstrated strong performance across domains, including text-to-image generation, inpainting, and video synthesis  [1, 24, 27, 12, 35, 39, 7]. More recently, the introduction of Diffusion Transformers (DiT) [23] has further advanced scalability and performance, enabling diffusion models to tackle complex generative scenarios and large-scale data [45, 20, 38, 48, 34, 40]. However, despite these advantages, the inherently iterative nature of the denoising process results in low inference speed. This limitation poses a significant challenge for deploying such models in low-resource or latency-sensitive applications. To address the inefficiency of diffusion inference, a variety of approaches have been proposed, reducing the number of sampling steps, such as designing advanced probabilistic-flow-inspired solvers based on Ordinary Differential Equations (ODEs) [17, 18, 44]. Other approaches focus on compressing model size by leveraging techniques like network pruning [46], knowledge distillation [11], and token reduction [9, 3, 41]. However, most of these methods require additional training or offline optimization, introducing extra computational costs and limiting the overall deployment flexibility.
@@
A recent line of work [2, 19, 28, 50, 49] explores the feature redundancy inherent in diffusion models. Based on the observation that features across adjacent steps often exhibit high similarity, these approaches aim to accelerate inference by reusing features from previous timesteps. By caching and reusing past representations instead of recomputing them, these methods significantly reduce computational overhead without any additional training effort. Nevertheless, as the gap between timesteps increases, the assumption of feature similarity gradually breaks down. Naively reusing cached features across distant steps can result in substantial errors, ultimately degrading image quality. To mitigate this, TaylorSeer [15] proposed to predict features at future timesteps using Taylor series expansion based on previously cached features, rather than directly reusing them. This enables the model to retain high generation quality, even under aggressive acceleration.
@@
While TaylorSeer improved the reliability of cached features by predicting future representations through Taylor expansion, it introduces considerable memory overhead and additional latency. Specifically, it applied Taylor expansion at every module within each transformer block (e.g., attention or feedforward module), predicting the residual branch in a step-by-step manner. However, modern diffusion transformer architectures consist of numerous modules, each requiring its own cached features. This design significantly increases memory consumption and computational burden, partially offsetting the intended acceleration gains. To address this issue, we propose a Last Block Forecast strategy that reduces memory usage with minimal impact on acceleration performance. TaylorSeer revealed that intermediate features across blocks are highly predictable. Given the strictly sequential nature of DiT, where each block passes information to the next, we propose to forecast only the last block’s output, which corresponds to the model output at the current timestep. By doing so, we obtain an accurate approximation of the last block output while avoiding the need to cache every module’s residual branch. This strategy requires storing only a single set of features and eliminates much of the computational overhead.
@@
Meanwhile, TaylorSeer adopted a fixed caching schedule by dividing the inference process into equally spaced intervals, computing one feature per interval and using Taylor expansion to predict the remaining ones. This design implicitly assumes that Taylor prediction remains consistently reliable across timesteps, which often does not hold. In practice, the model’s output exhibits dynamic changes, leading to fluctuations in approximation accuracy over time. As a result, a rigid reuse pattern may either miss opportunities for further acceleration or incur significant errors when prediction fails. This limitation raises a natural and important question: When is the Taylor predictor reliable enough to replace actual computation?
@@
Intuitively, we only want to rely on the Taylor predictor when it is sufficiently accurate; otherwise, we fall back to the standard full computation. However, evaluating the full computation output just to compare with the predicted result defeats the purpose of acceleration. To address this, we propose a Prediction Confidence Gating (PCG) mechanism. Our core hypothesis is that Transformer-based architectures like DiT exhibit strong sequential dependencies. In other words, the quality of early block predictions can reflect the predictability of later ones. Good predictions on early blocks suggest high reliability when forecasting the last block output. Based on this insight, we use the prediction error of the first block as a proxy to assess the confidence of the Taylor prediction at the current timestep. Concretely, we compute the first block’s true output at each timestep and apply Taylor expansion to predict it in parallel. If the predicted result closely matches the ground truth, we consider the model state to be sufficiently stable and apply Taylor approximation for the remaining computation. Otherwise, we fall back to full computation to avoid quality degradation. This approach maintains decision quality while incurring almost no additional computational cost, enabling a training-free, dynamic acceleration strategy.
@@ 2.1 Diffusion Transformer
Diffusion models [6, 29] have achieved remarkable success in content generation, producing high-quality and diverse outputs. Early approaches [24, 25, 27, 36] were predominantly based on U-Net architectures, which delivered strong performance in both image and video generation. However, their scalability is fundamentally limited, posing challenges for building large-scale, high-capacity models. To address this limitation, the Diffusion Transformer (DiT) [23] architecture was proposed, leveraging the scalability and modeling flexibility of Transformers. This shift has led to a wave of progress across various generative tasks. Nonetheless, the broader adoption of DiT remains constrained by the inherently iterative nature of the diffusion process, which significantly hampers inference efficiency.
@@ 2.2 Diffusion Model Acceleration
With the growing deployment of diffusion models in real-world applications, improving their inference efficiency has become an increasingly important research direction. A variety of methods have been proposed to accelerate the sampling process, which can generally be categorized into three main paradigms.
First, several works aim to shorten the sampling trajectory [32]. For example, methods like DDIM [30] reformulate the generative process to avoid long iterative updates, while others draw inspiration from numerical solvers to make the generation process more efficient. Second, model compression and quantization techniques have been explored to reduce computational overhead, including knowledge distillation [11], token reduction [9, 3, 41], and network pruning [46, 47]. However, these methods often require additional training or offline optimization, which introduces extra computational cost. Third, a growing body of work focuses on training-free acceleration via feature reuse [2, 19, 28, 50, 49], which has gained increasing attention due to its simplicity and effectiveness. These methods are motivated by the observation that intermediate features in diffusion models often contain redundancy, making it feasible to reuse them rather than recomputing at every step. For instance, PAB [43] and T-Gate [14] selectively reuse cached features based on different types of attention. TeaCache [13] proposes an adaptive caching strategy based on timestep-aware differences in output features. TaylorSeer [15] reduces the accumulation of cache reuse errors at distant timesteps by using Taylor expansion to predict feature values. Building upon TaylorSeer, our method further improves prediction utility by explicitly evaluating the reliability of the Taylor approximation at each step. This enables a more principled decision on the predictability of a timestep, leading to a better trade-off between generation quality and inference efficiency.
@@ 3 Method
Figure 2: The framework of our method. At each timestep $t$, we first compute the actual output of the first block and simultaneously predict it using Taylor expansion. The prediction error is then evaluated. If the error is below a threshold $\epsilon$, indicating that the Taylor prediction is reliable, we use it to approximate the last block feature-skipping the computation of the remaining B1B{-}1 blocks to accelerate inference. Otherwise, we fall back to full computation of all blocks. This process is repeated for each subsequent timestep.
@@ 3.1 Preliminary
TaylorSeer. Previous feature caching methods accelerate inference by reusing the same features across an interval of timesteps. Specifically, given an interval $\{t,t{-}1,\dots,t{-}N{+}1\}$, they compute $\mathcal{F}(\mathbf{x}_{t})$ once and approximate the rest as: $\mathcal{F}(\mathbf{x}_{t-k})\approx\mathcal{F}(\mathbf{x}_{t}),\text{for }k=1,\dots,N{-}1.$ This direct reuse strategy, while efficient, leads to increasing approximation error as the time gap widens. To address this, TaylorSeer [15] proposed to predict features rather than reuse them verbatim. Motivated by the observation that features and their derivatives evolve smoothly over timesteps, TaylorSeer applies Taylor series expansions using cached multi-order differences to capture the temporal dynamics of features. This method enables more accurate estimation of future features in a non-parametric manner. At timestep $t$, they defined a cache storing the feature and its $O$-th order finite differences:
@@ 3.2 Last Block Forecast
TaylorSeer revealed that intermediate features across blocks are highly predictable. Given the strictly sequential nature of DiT, where each block passes information to the next, we think that it is reasonable and feasible to accelerate the process by predicting the output of the last block. Based on this, we propose a Last Block Forecast strategy: instead of applying the Taylor predictor to every module’s intermediate features, we apply it directly to the last block at each timestep. Formally, given the last block feature $g(x_{t}^{B})$ at timestep $t$, we cache the following features and finite differences:
@@ 4.1 Experiment Settings
Model Configurations. To validate the generality of our method across diverse generative settings, we evaluate it on three representative diffusion models spanning different modalities and resolutions. FLUX.1-dev [10], a text-to-image model, produces high-resolution outputs of 1024×1024 using the Rectified Flow sampler [16]. DiT-XL/2 [23], a class-conditional image generator, follows a 50-step DDIM [30] schedule to generate 256×256 images. Meanwhile, Wan Video [33], designed for text-to-video generation, synthesizes 81-frame clips at 480×832 resolution, adhering to its original sampling configuration. This diverse selection of models ensures a comprehensive evaluation of both inference efficiency and generation quality.
Table 1: Quantitative comparison of our method and baselines on the FLUX model for text-to-image generation.
@@
FLUX.1 [10] FLOPs(T)↓ Latency(s)↓ Speed↑ Image Reward↑ SSIM ↑ PSNR↑
@@
PAB [43] 1169.60 14.80 1.57x 1.0006 0.8494 24.65
Teacache ($\delta$=0.25) [13] 1069.84 13.55 1.71x 1.0039 0.806 22.13
Taylorseer (NN=2, OO=1) [15] 993.53 13.56 1.71x 1.0005 0.850 24.02
@@
Teacache ($\delta$=0.40) [13] 802.48 10.14 2.29x 0.9889 0.716 19.26
Taylorseer (NN=5, OO=2) [15] 458.86 8.21 2.83x 0.9793 0.648 17.40
Ours ($\epsilon$=0.13) 503.72 7.19 3.13x 0.9956 0.813 22.00
@@
FLUX and Wan models are implemented in the Paddlepaddle [21] framework, while DiT is implemented using PyTorch [22]. To ensure fairness and compatibility, our acceleration method is integrated into each model’s original inference pipeline without altering their underlying architecture or training setup. All experiments are conducted on NVIDIA A800 80GB GPUs, and latency is measured under consistent hardware and software conditions.
Evaluation and Metrics. We evaluate the effectiveness of our method from two perspectives: acceleration and visual quality. Our experimental setup closely follows the configurations of TaylorSeer [15] and TeaCache [13] to ensure fair comparison. To assess inference efficiency, we uniformly report inference latency and Floating Point Operations (FLOPs) as the primary metric. For generation quality, we incorporate task-specific evaluation metrics tailored to each generation setting. For text-to-image generation, we adopt the DrawBench benchmark [27] and evaluate the results using ImageReward [37], which estimates alignment with human preferences. Furthermore, we use PSNR and SSIM to quantify the similarity between outputs produced by our accelerated method and those generated by the original models for text-to-image generation on the COCO-1K dataset. For class-conditional image generation, we employ the FID-50k [5], sFID-50k, and Inception Score, which measures the distributional similarity between generated samples and real training data. For text-to-video generation, we use VBench [8], a video-specific benchmark containing 946 prompts, designed to reflect human-aligned quality judgments. In addition to ImageReward, we further report PSNR, SSIM, and LPIPS [42] to assess the perceptual and structural similarity between videos generated by different methods.
@@
DiT [23] Attention [4] FLOPs(T)↓ Latency(s)↓ Speed↑ FID↓ sFID ↓ Inception Score↑
@@
FORA (N=3N=3[28] 8.58 0.197 2.17x 3.55 6.36 229.02
ToCa (N=3N=3[50] 10.23 0.216 1.98x 2.87 4.76 235.21
DuCa (N=3N=3[49] 9.58 0.208 2.06x 2.88 4.66 233.37
Taylorseer (N=3,O=3N=3,O=3[15] 8.56 0.248 1.73x 2.34 4.69 238.42

### v3

@@ 1 Introduction
Figure 2: The expanded architecture of the DiT and the difference between different methods. (a) DiTs follow a hierarchical architecture which are composed of multiple transformer blocks, where each block typically consists of three modules. (b) In the original method, every module is computed at each timestep, resulting in a high computational cost. TaylorSeer reduced computation by calculating a module at one timestep and using Taylor expansions to predict its features for the remaining timesteps within a given interval NN. Our method uses a confidence-gated method: if the first block’s prediction error at timestep t-1 is small, we directly predict the last block’s feature at this timestep, skipping computation of the B1B-1 blocks to achieve acceleration. Otherwise, we compute all blocks normally.
@@
In recent years, diffusion models  [ho2020denoising, sohl2015deep, song2019generative, ke2024text, liu2025diffosr, qu2024mtlsc] have become foundational to visual generation, owing to their impressive ability to produce high-quality and semantically rich images. Initially built upon U-Net [ronneberger2015u] architectures, these models have demonstrated strong performance across domains, including text-to-image generation, inpainting, and video synthesis  [blattmann2023stable, ramesh2022hierarchical, li2025denoising, wang2025cymodiff, yuan2025diff, hu2025dtiu]. More recently, the introduction of Diffusion Transformers (DiT) [peebles2023scalable] has further advanced scalability and performance, enabling diffusion models to tackle complex generative scenarios and large-scale data [zheng2024open, yang2024cogvideox, zhuang2025enhanced, wang2025diff, zeng2025enhancing]. However, despite these advantages, the inherently iterative nature of the denoising process results in low inference speed. This limitation poses a significant challenge for deploying such models in low-resource or latency-sensitive applications. To address the inefficiency of diffusion inference, a variety of approaches have been proposed, reducing the number of sampling steps, such as designing advanced probabilistic-flow-inspired solvers based on Ordinary Differential Equations (ODEs) [lu2022dpm, lu2025dpm, zheng2023dpm]. Other approaches focus on compressing model size by leveraging techniques like network pruning [zhu2024dip], knowledge distillation [li2023snapfusion], and token reduction [kim2024token, cheng2025cat]. However, most of these methods require additional training or offline optimization, introducing extra computational costs and limiting the overall deployment flexibility.
@@
A recent line of work [chen2024delta, ma2024deepcache, selvaraju2024fora, zou2024accelerating, zou2024acceleratingduca] explores the feature redundancy inherent in diffusion models. Based on the observation that features across adjacent steps often exhibit high similarity, these approaches aim to accelerate inference by reusing features from previous timesteps. Diffusion Transformers (DiTs) follow a hierarchical architecture as illustrated in Figure 2. They are composed of multiple transformer blocks, where each block typically consists of three modules: cross-attention, self-attention, and feed-forward networks. During inference, every timestep must pass through all these modules, leading to substantial computational overhead. Intuitively, if the model could reuse the intermediate representations computed in earlier timesteps, rather than recalculating them from scratch at each timestep, the inference process would become significantly more efficient. By caching and reusing past representations instead of recomputing them, these methods reduce redundant computation and achieve faster sampling without requiring any additional training effort. Nevertheless, as the gap between timesteps increases, the assumption of feature similarity gradually breaks down. Naively reusing cached features across distant steps can result in substantial errors, ultimately degrading image quality. To mitigate this, TaylorSeer [liu2025reusing] proposed to predict features at future timesteps using Taylor series expansion based on previously cached features, rather than directly reusing them. This enables the model to retain high generation quality, even under aggressive acceleration settings.
@@
While TaylorSeer improved the reliability of cached features by predicting future representations through Taylor expansion, it introduces considerable memory overhead and additional latency. Specifically, it applied Taylor expansion at every module within each transformer block (e.g., attention or feedforward module) as shown in Figure 2, predicting the residual branch in a step-by-step manner. However, modern diffusion transformer architectures consist of numerous modules, each requiring its own cached features. This design significantly increases memory consumption and computational burden, partially offsetting the intended acceleration gains. To address this issue, we propose a Last Block Forecast strategy that reduces memory usage with minimal impact on acceleration performance. TaylorSeer revealed that intermediate features across blocks are highly predictable. Given the strictly sequential nature of DiT, where each block passes information to the next, we propose to forecast only the last block’s output, which corresponds to the model output at the current timestep. By doing so, we bypass the computations of the intermediate blocks to accelerate inference, while still avoiding the need to cache every module’s residual branch. This strategy modifies TaylorSeer’s original acceleration mechanism of forecasting instead of reuse, and it requires storing only a single set of features, thereby eliminating much of the computational overhead.
@@
Meanwhile, TaylorSeer adopted a fixed caching schedule by dividing the inference process into equally spaced intervals. As shown in Figure 2, given an interval t,t1,,tN+1{t,t-1,...,t-N+1}, TaylorSeer will compute the feature at t and predict the rest of the timesteps using Taylor expansion. This design implicitly assumes that Taylor prediction remains consistently reliable across timesteps, which often does not hold. In practice, the model’s output exhibits dynamic changes, leading to fluctuations in approximation accuracy over time. As a result, a rigid reuse pattern may either miss opportunities for further acceleration or incur significant errors when prediction fails. This limitation raises a natural and important question: When is the Taylor predictor reliable enough to replace actual computation?
@@
Intuitively, we only want to rely on the Taylor predictor when it is sufficiently accurate; otherwise, we fall back to the standard full computation. However, evaluating the full computation output just to compare with the predicted result defeats the purpose of acceleration. To address this, we propose a Prediction Confidence Gating (PCG) mechanism. Our core hypothesis is that Transformer-based architectures like DiT exhibit strong sequential dependencies. In other words, the quality of early block predictions can reflect the predictability of later ones. Good predictions on early blocks suggest high reliability when forecasting the last block output. Based on this insight, we use the prediction error of the first block as a proxy to assess the confidence of the Taylor prediction at the current timestep. Concretely, we compute the first block’s true output at each timestep and apply Taylor expansion to predict it in parallel. If the prediction error with respect to the ground truth is sufficiently small, we consider the model state to be sufficiently stable and apply the Taylor approximation for the remaining computation. Otherwise, we fall back to full computation to avoid quality degradation. This approach maintains decision quality while incurring almost no additional computational cost, enabling a training-free, dynamic acceleration strategy.
@@ 2.1 Diffusion Transformer
With the rapid development of deep learning [zhao2024mixir, liu2023reliable, qian2025visible], diffusion models [ho2020denoising, sohl2015deep, liu2024iterative, wang2024silent, huang2025implicit] have achieved remarkable success in content generation, producing high-quality and diverse outputs. Early approaches [ramesh2022hierarchical, rombach2022high] were predominantly based on U-Net architectures, which delivered strong performance in both image and video generation. However, their scalability is fundamentally limited, posing challenges for building large-scale, high-capacity models. To address this limitation, the Diffusion Transformer (DiT) [peebles2023scalable] architecture was proposed, leveraging the scalability and modeling flexibility of Transformers. This shift has led to a wave of progress across various generative tasks. Nonetheless, the broader adoption of DiT remains constrained by the inherently iterative nature of the diffusion process, which significantly hampers inference efficiency.
@@ 2.2 Diffusion Model Acceleration
With the growing deployment of diffusion models in real-world applications, improving their inference efficiency has become an increasingly important research direction. A variety of methods have been proposed to accelerate the sampling process, which can generally be categorized into four main paradigms.
First, several works aim to shorten the sampling trajectory [song2023consistency]. For example, methods like DDIM [song2020denoising] reformulate the generative process to avoid long iterative updates, while others draw inspiration from numerical solvers to make the generation process more efficient. Second, model compression and quantization techniques have been explored to reduce computational overhead, including knowledge distillation [li2023snapfusion], token reduction [kim2024token, cheng2025cat], and network pruning [zhu2024dip, zhu2025dtsformer]. Third, a set of training-oriented architecture modification methods aims to improve efficiency by altering the model design during training. For example, the Transformer architecture can be adapted to Mamba architecture [huang2025m4v], attention modules can be replaced with Mamba attention [gao2024matten], new modules can be introduced within the Transformer structure [po2025long], or the Transformer can be designed without residual prediction [jiao2025flexvar] to reduce inference cost. However, these methods often require additional training or offline optimization, which introduces extra computational cost. Fourth, a growing body of work focuses on training-free acceleration via feature reuse [chen2024delta, ma2024deepcache, selvaraju2024fora, zou2024accelerating, zou2024acceleratingduca], which has gained increasing attention due to its simplicity and effectiveness. These methods are motivated by the observation that intermediate features in diffusion models often contain redundancy, making it feasible to reuse them rather than recomputing at every step. For instance, PAB [zhao2024real] and T-Gate [liu2024faster] selectively reuse cached features based on different types of attention. TeaCache [liu2025timestep] proposes an adaptive caching strategy based on timestep-aware differences in output features. TaylorSeer [liu2025reusing] reduces the accumulation of cache reuse errors at distant timesteps by using Taylor expansion to predict feature values. Building upon TaylorSeer, our method further improves prediction utility by explicitly evaluating the reliability of the Taylor approximation at each step. This enables a more principled decision on the predictability of a timestep, leading to a better trade-off between generation quality and inference efficiency.
@@ 3.1 Preliminary
Figure 3: The framework of our method. At timestep t1t-1, we first compute the actual output of the first block and simultaneously predict it using Taylor expansion with the cached feature at timestep $t$. The prediction error is then evaluated. If the error is below a threshold $\epsilon$, the Taylor prediction is considered reliable. In this case, we use it to approximate the last block feature, thereby skipping the computation of the remaining (B1)(B-1) blocks to accelerate inference. Otherwise, we fall back to full computation of all blocks. This process is repeated for each subsequent timestep.
@@
TaylorSeer. Previous feature caching methods accelerate inference by reusing the same features across an interval of timesteps. Specifically, given an interval $\{t,t{-}1,\dots,t{-}N{+}1\}$, they compute $\mathcal{F}(\mathbf{x}_{t})$ once and approximate the rest as: $\mathcal{F}(\mathbf{x}_{t-k})\approx\mathcal{F}(\mathbf{x}_{t}),\text{for }k=1,\dots,N{-}1.$ This direct reuse strategy, while efficient, leads to increasing approximation error as the time gap widens. To address this, TaylorSeer [liu2025reusing] proposed to predict features rather than reuse them verbatim. Motivated by the observation that features and their derivatives evolve smoothly over timesteps, TaylorSeer applies Taylor series expansions using cached multi-order differences to capture the temporal dynamics of features. This method enables more accurate estimation of future features in a non-parametric manner. At timestep $t$, they defined a cache storing the feature and its $O$-th order finite differences:
@@ 3.2 Last Block Forecast
TaylorSeer revealed that intermediate features across blocks are highly predictable. Given the strictly sequential nature of DiT, where each block passes information to the next, we think that it is reasonable and feasible to accelerate the process by predicting the output of the last block. In this way, we can bypass the computations of the intermediate blocks while still obtaining the DiT output at the current timestep. Based on this, we propose a Last Block Forecast strategy: instead of applying the Taylor predictor to every module’s intermediate features, we apply it directly to the last block at each timestep. Formally, given the last block feature $g(x_{t}^{B})$ at timestep $t$, we cache the following features and finite differences:
@@ 4.1 Experiment Settings
Table 1: Quantitative comparison of our method and baselines on FLUX for text-to-image generation.
@@
FLUX.1 [flux] FLOPs(T)↓ Latency(s)↓ Speed↑ Image Reward↑ SSIM ↑ PSNR↑
@@
PAB [zhao2024real] 1169.60 14.80 1.57x 1.0006 0.849 24.65
Teacache ($\delta=0.25$[liu2025timestep] 1069.84 13.55 1.71x 1.0039 0.806 22.13
Taylorseer ($N$ = 2, $O$ = 1) [liu2025reusing] 993.53 13.56 1.71x 1.0005 0.850 24.02
@@
Teacache ($\delta$ = 0.40) [liu2025timestep] 802.48 10.14 2.29x 0.9889 0.716 19.26
Taylorseer ($N$ = 5, $O$ = 2) [liu2025reusing] 458.86 8.21 2.83x 0.9793 0.648 17.40
Ours ($\epsilon$ = 0.13) 503.72 7.19 3.13x 1.0216 0.813 22.00
Table 2: Quantitative comparison of our method and baselines on the FLUX model for text-to-image generation on the GenEval benchmark.
Method Acceleration GenEval
FLUX.1 [flux] Speed↑ overall↑ Single Obj.↑ Two Obj.↑ Counting↑ Colors↑ Position↑ Attr. Binding↑
dev 50steps 1.00x 0.67 1.00 0.83 0.73 0.78 0.21 0.48
PAB [zhao2024real] 1.57x 0.66 0.99 0.82 0.72 0.76 0.21 0.47
Teacache ($\delta$ = 0.25) [liu2025timestep] 1.71x 0.66 0.99 0.81 0.73 0.78 0.18 0.46
Taylorseer(N=2,O=1N=2,O=1[liu2025reusing] 1.71x 0.67 0.99 0.83 0.73 0.79 0.20 0.49
Ours($\epsilon$ = 0.03) 2.06x 0.67 1.00 0.84 0.73 0.77 0.20 0.48
Teacache ($\delta$ = 0.40) [liu2025timestep] 2.29x 0.67 1.00 0.83 0.72 0.79 0.19 0.47
Taylorseer(N=5,O=2N=5,O=2[liu2025reusing] 2.83x 0.65 0.99 0.84 0.71 0.45 0.21 0.45
Ours($\epsilon$ = 0.13) 3.13x 0.67 0.99 0.81 0.77 0.77 0.21 0.48
Model Configurations. To validate the generality of our method across diverse generative settings, we evaluate it on four representative diffusion models spanning different modalities and resolutions. FLUX.1-dev [flux], a text-to-image model, produces high-resolution outputs of 1024×1024 using the Rectified Flow sampler [liu2022flow]. DiT-XL/2 [peebles2023scalable], a class-conditional image generator, follows a 50-step DDIM [song2020denoising] schedule to generate 256×256 images. Meanwhile, Wan Video [wan2025wan], designed for text-to-video generation, synthesizes 81-frame clips at 480×832 resolution, adhering to its original sampling configuration. Hunyuan Video [kong2024hunyuanvideo] also designed for text-to-video generation, generates 61-frame clips at 320×512 resolution using the default diffusers implementation, with the only modification being that we set the inference steps to 50. This diverse selection of models ensures a comprehensive evaluation of both inference efficiency and generation quality.
FLUX, Wan Video, and Hunyuan Video models are implemented in the PaddlePaddle [ma2019paddlepaddle] framework, while DiT is implemented using PyTorch [paszke2019pytorch]. To ensure fairness and compatibility, our acceleration method is integrated into each model’s original inference pipeline without altering its underlying architecture or training setup. All experiments are conducted on NVIDIA A800 80GB GPUs, and latency is measured under consistent hardware and software conditions.
Evaluation and Metrics. We evaluate the effectiveness of our method from two perspectives: acceleration and visual quality. Our experimental setup closely follows the configurations of TaylorSeer [liu2025reusing] and TeaCache [liu2025timestep] to ensure fair comparison. To assess inference efficiency, we uniformly report inference latency and Floating-Point Operations (FLOPs) as the primary metrics. For generation quality, we incorporate task-specific evaluation metrics tailored to each generation setting. For text-to-image generation, we adopt the DrawBench benchmark [saharia2022photorealistic] and evaluate the results using ImageReward [xu2023imagereward], which estimates alignment with human preferences. In addition, we employ GenEval [ghosh2023geneval], an object-focused framework to evaluate compositional image properties across multiple dimensions. Furthermore, we use PSNR and SSIM to quantify the similarity between outputs produced by our accelerated method and those generated by the original models for text-to-image generation on the COCO-1K dataset. For class-conditional image generation, we employ the FID-50k [heusel2017gans], sFID-50k, and Inception Score, which measures the distributional similarity between generated samples and real training data. For text-to-video generation, we use VBench [huang2024vbench], a video-specific benchmark containing 946 prompts, designed to reflect human-aligned quality judgments. In addition to Vbench, we further report PSNR, SSIM, and LPIPS [zhang2018unreasonable] to assess the perceptual and structural similarity between videos generated by different methods.
@@
DiT [peebles2023scalable] Attention [dao2022flashattention] FLOPs(T)↓ Latency(s)↓ Speed↑ FID↓ sFID ↓ Inception Score↑
@@
FORA (N=3N=3[selvaraju2024fora] 8.58 0.197 2.17x 3.55 6.36 229.02
ToCa (N=3N=3[zou2024accelerating] 10.23 0.216 1.98x 2.87 4.76 235.21
DuCa (N=3N=3[zou2024acceleratingduca] 9.58 0.208 2.06x 2.88 4.66 233.37
Taylorseer (N=3,O=3N=3,O=3[liu2025reusing] 8.56 0.248 1.73x 2.34 4.69 238.42
@@
Table 4: Quantitative comparison of our method and baselines on Wan Video for text-to-video generation.
Method Acceleration Visual Quality
Wan Video [wan2025wan] Latency(s)↓ Speed↑ VBench(%)↑ SSIM ↑ PSNR↑ LPIPS ↓
ori 50steps 165.69 1.00x 82.72 - - -
PAB [zhao2024real] 78.02 2.12x 73.60 0.5205 17.1515 0.4783
Teacache ($\delta=0.26$[liu2025timestep] 43.25 3.83x 76.67 0.4147 12.4317 0.6332
Taylorseer (N=5,O=1N=5,O=1[liu2025reusing] 45.27 3.66x 78.71 0.4167 12.6739 0.5758
Ours ($\epsilon=0.36$) 40.02 4.14x 81.86 0.6319 17.2108 0.2882
Table 5: Quantitative comparison of our method and baselines on Hunyuan Video for text-to-video generation.
Method Acceleration Visual Quality
Hunyuan Video [kong2024hunyuanvideo] Latency(s)↓ Speed↑ VBench(%)↑ SSIM ↑ PSNR↑ LPIPS ↓
ori 50steps 101.48 1.00x 80.18 - - -
PAB [zhao2024real] 47.35 2.14x 79.16 0.6193 16.8222 0.3707
Teacache ($\delta=0.15$[liu2025timestep] 40.55 2.50x 79.71 0.6602 18.1942 0.3049
Taylorseer (N=5,O=1N=5,O=1[liu2025reusing] 30.20 3.36x 78.88 0.5765 15.4932 0.4309
Ours ($\epsilon=0.12$) 27.45 3.70x 79.75 0.7287 20.5806 0.2159